A partial adjusted trial balance of Swifty Company at January 31, 2020, shows the following.

SWIFTY COMPANY
ADJUSTED TRIAL BALANCE
JANUARY 31, 2020

Debit

Credit

Supplies $820
Prepaid Insurance 3,120
Salaries and Wages Payable $920
Unearned Service Revenue 870
Supplies Expense 950
Insurance Expense 520
Salaries and Wages Expense 1,920
Service Revenue 2,120

Answer the following questions, assuming the year begins January 1.
If the amount in Supplies Expense is the January 31 adjusting entry, and $850 of supplies was purchased in January, what was the balance in Supplies on January 1?
Beginning balance of supplies $

If the amount in Insurance Expense is the January 31 adjusting entry, and the original insurance premium was for one year, what was the total premium?
Total premium $


When was the policy purchased?
The policy was purchased on

July 1, 2019August 1, 2019August 31, 2019

.
If $2,620 of salaries was paid in January, what was the balance in Salaries and Wages Payable at December 31, 2019?
Beginning balance of salaries and wages payable $

If $1,720 was received in January for services performed in January, what was the balance in Unearned Service Revenue at December 31, 2019? Assume that there are no accounts receivable.
Beginning balance of unearned service revenue $
Answer #1

1) Beginning balance of supplies = 950+820-850 = 920

2) Total premium = 520*12 = 6240

Premium expired = (6240-520-3120) = 2600

So 5 month insurance expired before Jan 1

So answer is August 1, 2019

3) Beginning balance of salaries and wages payable = 2620+920-1920 = 1620

4) Beginning balance of unearned service revenue = 1720+870 = 2590